if the square of a positive integer is added to three times the integer is added to three times the integer, the result is 28. find the integer.
Answer by ankor@dixie-net.com(22740)

If the square of a positive integer is added to three times the integer, the result is 28.
find the integer.
:
Write an equation for what it says
x^2 + 3x = 28
a quadratic equation
x^2 + 3x - 28 = 0
factors to:
(x+7)(x-4) = 0
the positive solution
x = 4
:
:
See if that checks out
4^2 + 3(4) = 28
16 + 12 = 28